27 February 1996
Special acquisitions have been planned for the period 10 February to 1 March 1996 due to the Baltic Sea icing over.
20 February 1996
A special tandem acquisition related to the oil spill off the coast of Milford Haven, Wales (UK), has been planned for 25-26 February.
16 February 1996
The comparison of GOME level 1 and level 2 products with data measured by other instruments gave satisfactory results. For the first time minor trace gases highly influencing the ozone depletion can be measured from space. Product distribution may start in April 1996. More details and Preliminary Antartic Ozone Map in GOME report.
14 February 1996
January 1996 update of ERS-1 SAR Stability Monitoring.
05 February 1996
User proposals for the new announcement of opportunity being made by ESA to allow further interferometry scientific studies between ERS-1 and ERS-2 can be made up to 15th March 1996. For applications, please refer to your National ESA delegation. For further information please contact the ERS Helpdesk at ESA-ESRIN, (Italy).
31 January 1996
Due to an increase in detector temperature and in order to maximise the remaining life time of the ATSR-1 instrument, ATSR-1 is operating nominally in 10 bits mode since 12 January 1996, and should continue to do so until the end of the mission. This is a new mode of in-flight operation for ATSR-1 which should be handled correctly in the ground processing software, but which has not been tested extensively with flight data. Because the signal to noise performance of ATSR-1 was already degrading, initial indications show that the truncation of the signal channel telemetry to 10 bits has little effect on the quality of the spatially averaged data set. RAL (Rutherford Appleton Laboratory) are in process of re-validating the ASST products and a further announcement on data quality will be made during the next month.
31 January 1996
The AMI/SAR data - available to users since July 1995 - keep being distributed. Acquisition is tailored according to the requirements of interferometric studies. The interferometry campaign has been fixed so far to last unitl 17th May 1996. A new announcement of opportunity is being made by ESA to allow further exploitation of ERS-1 and ERS-2 tandem mission. Closing date is 15th March 1996 for the user proposal. For further information please refer to your National ESA delegation or to the ERS Helpdesk.
30 January 1996
The following acquisitions, for both ERS-1 and ERS-2, will be planned over the flooded area in southern France (29 Jan. 1996).
23 January 1996
The Along Track Scanning Radiometer (ATSR-2) which was working remarkably well since launch, ceased normal operations on 22 December 1995 because of a problem with the scanning mirror. A team at the Central Laboratories of the Research Council (CLRC) in the United Kingdom, which developed and funded this UK-provided instrument, is investigating the reasons for the anomaly. CLRS has agreed to keep the media informed of progress in the investigation and of the efforts made to resume the instrument's normal operations.
Showing 4,991 - 5,000 of 5,102 results.